OpenStudy (anonymous):

What is the solution to the rational equation 5 over x squared plus 3x plus 2 + 1 over x plus 2 = 1 over 3x plus 3

OpenStudy (e.mccormick):

\(\dfrac{5}{x^2+3x+2}+\dfrac{1}{x+2}=\dfrac{1}{3+3x}\) ??

OpenStudy (anonymous):



OpenStudy (e.mccormick):

Factor it, find an LCD, add up all the fractions, then see what you get.

OpenStudy (anonymous):

i forgot i had to factor it

OpenStudy (anonymous):

thanks i got it from here

OpenStudy (e.mccormick):

Ah, Kk. Yah, for the LCD. \(\dfrac{5}{(x+1)(x+2)}+\dfrac{1}{x+2}=\dfrac{1}{3(x+1)}\) The nice thing about the simple ones is the other terms give clues on the factoring.